OpenStudy (anonymous):

the ratio of the measure of two complementary angles is 4 to 1. find the measure of each angle. help??

OpenStudy (anonymous):

Okie dokie so Complementary angles add to 180 4x+x=180 5x=180 x=36 so the angles are 36 and 4(36) which is 144 144 +36 180! =)

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (anonymous):

Woops!! Sorry!! I made a mistake. Complementary angles add up to 90, so just follow the steps but add them up to 90; supplementary angles are the ones that add up to 180. =) My mistake!

OpenStudy (anonymous):

@sarahnorville and you are sooooo welcome!!

OpenStudy (anonymous):

oh no problem! so itd be 5x= 90, x= 18 and 4(18)?

OpenStudy (anonymous):

Yes! Good job. =)

OpenStudy (anonymous):

awesome!:D thanks again!

OpenStudy (anonymous):

My pleasure! If you need anything else, I'm here! =)

OpenStudy (anonymous):

okay, so im having trouble with this one: "A wire 60 inches long is cut into two pieces whose lengths have a ratio of 3 to 1. To find the length of each piece, let x be the length of the shorter piece. Since the wire is 60 inches long, and since the shorter piece is x inches long, what is the length of the longer piece?" how do i start that problem exactly?

OpenStudy (anonymous):

would it just be 4x= 60, x=15 and 45?

OpenStudy (anonymous):

@catsarecool1995

OpenStudy (anonymous):

All righty! If the ration of the lengths is 3 to 1, the ratios of the lengths to the total length is 3 to 4 and 1 to 4. Given: total length = 60 inches shorter piece= x in \[1/4 = x/60\] multiply both sides by 60 15=x Since the total length is 60, the longer piece is 60-15=45 in Final answer: 45 inches!

OpenStudy (anonymous):

Hope that helps clear it up for you!

OpenStudy (anonymous):

okay, yeah that makes sense!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Yay! Glad it makes sense. Your welcome! =)

OpenStudy (anonymous):

@catsarecool1995 are you good at multiplying rational exponents by chance??

OpenStudy (anonymous):

Let me try! What is your question?

OpenStudy (anonymous):

well this is "equations with radicals" and its the square root of 2x+5 equals 7

OpenStudy (anonymous):